OpenStudy (anonymous):

>_< Help!! solve the equation 155 = -4x - 5(4x - 7)

OpenStudy (anonymous):

?

OpenStudy (epikrebel):

You start off by multiplying the parentheses. \[155=-4x-(5*4x)-(5*-7)\] \[155=-4x-20x+35\] \[155(-35)=-24x+35(-35)\] \[\frac{ 120 }{ -24 }=\frac{-24x}{-24}\] You get x=-5

OpenStudy (epikrebel):

Do you understand?

OpenStudy (anonymous):

Oh i seee. ;) thank you very much.! that helped alot